Overhead costs are overapplied if the amount applied to Work in Process is: Multiple Choice

greater than estimated overhead.

less than estimated overhead.

greater than actual overhead incurred.

less than actual overhead incurred.
Answer #1

Overhead costs are overapplied if the amount applied to Work in Process is

Answer- greater than actual overhead incurred

Example: Manufacturing overhead was estimated to be $4,00,000 along with 20,000 direct labour hours. Actual manufacturing overhead was $4,15,000 and actual labour hours were 21,000 hours.

Example answer- Predetermined overhead rate= $4,00,000÷20,000= $20.00

Actual manufacturing overhead= $4,15,000÷21,000= $20×21,000= $4,20,000

Overapplied overhead=$4,20,000-$4,15,000= $5,000